OpenStudy (anonymous):

how to solve 1 5 - x 3 3

hero (hero):

\[5 \frac{1}{3} + 5 \frac{1}{3} + 5\frac{1}{3} = (5 + 5 + 5) + \left(\frac{1}{3} + \frac{1}{3} + \frac{1}{3}\right) = 15 + \frac{3}{3} = 16\]

OpenStudy (anonymous):

thanks

hero (hero):

Or, alternatively, you can just do it the usual way, by changing the mixed fraction to an improper one: \[\frac{16}{3} \times 3 = 16\]